OpenStudy (anonymous):

Solve the system of linear equations below. 4x - 3y = 17 5x + 4y = 60

OpenStudy (anonymous):

(4x - 3y = 17) *4 (5x + 4y = 60 ) *3

OpenStudy (anonymous):

16x - 12y = 68 15x + 12y = 180 x = 248 y=325